OpenStudy (anonymous):

What is the value of the expression –9 – 7 + 5? A. 11 B. –11 C. 3 D. –7

OpenStudy (anonymous):

(I think b)

OpenStudy (texaschic101):

-9 - 7 + 5 = - 16 + 5 = - 11.....yep, you got it

OpenStudy (anonymous):

texaschic is right.

OpenStudy (texaschic101):

:) confirmation is always good